Two plane mirrors are inclined to each other such that a ray incident to the first mirror (M1) and parralel to second mirror (M2) is finally reflected from the second mirror (M2) parralel to the first mirror (M1). The angle between the two mirrors is?
  1. 45 degrees
  2. 60 degrees
  3. 75 degrees
  4. 90 degrees
Plz explain...

Grade:12th pass

2 Answers

Arun
3 years ago
Law of reflection -
 
1) Incident ray, reflected rays and normal at the point of incidence lie in the same plane.
 
2) Angle of incidence is equal to angle of reflection.
 
- wherein
 
L_{i}= L_{r}
 
x+2(90-\theta )+\theta =180

x=\theta
 
y+2(90-\theta )+\theta =180
 
=>y=\theta
 
from \: \: \Delta ABC
 
3\theta =180
 
=>\theta =60^{\circ}
 
Ayush_Deep
3 years ago
Dear friend the correct answer is 60° . The explanation is little bit tough without ray diagram.
Let assume the angle between the mirror is ¢.
The incident ray is parallel to one mirror and the final reflected ray is parallel to othe rmirror 
So ,angle between the rays will be 180-¢ including direction of ray . Angle of deviations is 180-¢
As we know that the angle of deviation is also equal to 2¢ so
180-¢ = 2¢
And ¢=60°
Hope you understood my method . This is also a correct method.